# The Population Resemblance Statistic: A Chi-Square Measure of Fit for Banking

### v3

@@ Abstract
Keywords: credit model risk; discrete goodness-of-fit; non-central chi-square; population stability index (PSI); model validation and monitoring; Kullback-Leibler divergence.
@@ 1 Introduction
Testing the stability of a population used for model development is common practice in model risk management. In credit risk modeling, the Population Stability Index (PSI) is the most widely used measure to monitor the evolution of the population underlying a model, through assessing the degree of discrepancy, conversely similarity, between two discrete probability distributions (see Thomas et al., (2002, pp. 155 ff.) and Siddiqi, (2017, pp. 368 ff.)). Small deviations in the population can result in inaccurate or unreliable model predictions. For example, consider modeling the Probability of Default (PD), or risk score, based on a given population of borrowers; if the latter changes substantively, the reliability of the PD model becomes questionable. In this case, the current population should resemble the one used during model development as it is a requirement of the prudential authorities when the model is used in the calculation of regulatory capital (see European Central Bank, (2024, 2019), Board of Governors of the Federal Reserve System, (2011), South African Reserve Bank, (2022); and Pruitt, (2010) for an application of the PSI in $SAS^{\text{\textregistered}}$). The same holds true for modeling Exposure at Default (EAD) or Loss Given Default (LGD) in credit risk. Other areas of application include insurance, healthcare, engineering, and marketing (see Huang et al., (2022), Li et al., (2022), Sahu et al., (2023), Dong et al., (2022), Wu and Olson, (2010), McAdams et al., (2022), Chou et al., (2022), Karakoulas, (2004) and Brockett et al., (1995)).
@@
Despite its widespread use, the origins and properties of the PSI are not widely understood. The PSI is based on the Kullback-Leibler divergence, measuring difference between two probability distributions (Kullback and Leibler, , 1951, eq. (2.6)). The earliest reference to the PSI measure can be found in Lewis, (1994), who also coined the term “Population Stability Index” and popularized use of the so-called Lewis constants as thresholds (see Thomas et al., (2002, p. 155 ff.), Siddiqi, (2017, p. 368 ff.)). Lewis, (1994, p. 106) describes the PSI without formulating a hypothesis in the statistical sense but notes, “If a user finds the distribution of scores close together, [they] can be confident that the population has not changed.” In his example, a PSI below 0.10 indicates that the current population resembles the original and no action is required, a value between 0.10 and 0.25 suggests that some investigation should be undertaken, and a value above 0.25 signals a substantial change in the incoming population that may necessitate model reconstruction. In our paper, the word resembles is used in the sense of difference by no more than a specified small deviation – see Definition 1 in Section 2.
@@
The arbitrary nature of the Lewis constants has been acknowledged by authors and practitioners alike (Yurdakul and Naranjo, (2020), Du Pisanie and Visagie, (2020) and Peters, (2021)). These thresholds pose significant limitations in portfolios with a small number of borrowers, where the shift of a single borrower results in a distortion of the PSI beyond its thresholds, thereby unnecessarily prompting model reconstruction. See Nedbank Group, (2023, p. 71), Standard Bank Group, (2024, p. 51) and FirstRand, (2024, p. 232) for the prevalence of portfolios with less than (e.g.) 100 borrowers. Conversely, in large portfolios (e.g. $>1$ million borrowers in a retail portfolio), shifts of large volumes of borrowers may remain undetected. These scenarios highlight the PSI’s limitations as a universal discrepancy measure and emphasize the need for practitioners to interpret the results with caution.
@@ 1.1 Research aims and objectives
This paper introduces the Population Resemblance Statistic (PRS) as a novel and easy-to-use alternative to the PSI for population resemblance monitoring. Based on the Pearson chi-square statistic, the PRS leverages the non-central chi-square distribution to derive critical values adjusted for sample size and number of risk categories. This approach allows for a more nuanced detection of population shifts, particularly in sample settings where the PSI is known to produce unreliable results. Furthermore, this paper lists some additional measures from the literature to provide greater context for population monitoring. The key objectives of this research are as follows:
@@
  • To introduce the concept of resemblance as a convenient way of constraining population shift, which is easily communicable to practitioners and that can be utilized to derive business outcomes.
  • To develop and formalize the PRS as an alternative population monitoring measure that addresses the identified limitations of the PSI, through the setting of well-founded critical values for action.
  • To formulate the PRS in a setting of a composite null hypothesis where a point null hypothesis would be too strict, giving rise to use of the least favorable non-central chi-square distribution for decision making.
  • To evaluate the performance of the PRS through simulation studies and real-world applications, highlighting its competency in both small and large sample sizes.
@@
1.2 Paper structure
The remainder of this paper is structured as follows: Section 2 formalizes the problem, introducing the PRS as an alternative discrepancy measure with a brief discussion of related methods. Section 3 outlines the statistical properties of the PRS, while Section 3.3 details the derivation of sample-size dependent decision boundaries (critical values). Section 3.4 contains a practical guide to implementing the PRS, also giving a flowchart for its application. Section 4 presents a comprehensive simulation study, and Section 5 applies the PRS to real-world data, comparing it with the widely used PSI and other measures. Finally, Section 6 summarizes key contributions and suggests directions for future research.
@@ 2.1 Statistical representation
Consider an independent and identically distributed (i.i.d.) sample of ordinal scores, $X_{1},X_{2},\ldots,X_{n}$ from a discrete population with cumulative distribution function (cdf) $F$ defined on the set of integers $\{1,2,\ldots,B\}$. These ordinal scores often arise by discretizing continuous or numerical values into predefined categories, with each score representing membership in one of $B\geq 2$ disjoint categories, reflecting, for example, level of risk.
Let $n_{i}$ denote the count of scores in category $i$, formally expressed as $n_{i}=\sum_{j=1}^{n}\mathbb{I}(X_{j}=i)$, where $\mathbb{I}(A)$ is the indicator function such that $\mathbb{I}(A)=1$ when $A$ is true and $\mathbb{I}(A)=0$ otherwise. The total number of observed scores, $n$, is given by the sum of all category counts, $n=\sum_{i=1}^{B}n_{i}$.
The true probability of a score falling into category $i$ is $p_{i}=P(X_{j}=i)=F(i)-F(i-1)$ for $i=1,\ldots,B$. Let $\bm{p}=(p_{1},p_{2},\ldots,p_{B})^{\top}$. The observed category proportions, serving as unbiased estimators of the true probabilities, are denoted by $\widehat{\bm{p}}=(\hat{p}_{1},\hat{p}_{2},\ldots,\hat{p}_{B})^{\top}$ with each $\hat{p}_{i}=n_{i}/n$ for $i=1,2,\ldots,B$.
The aim is to determine whether the current population $\bm{p}$ resembles the reference population $\bm{p}_{0}=(p_{01},p_{02},\ldots,p_{0B})^{\top}$, which was used to construct the model. The specific nature of this model is not germane. Rather, the question is whether $\bm{p}$ has “substantively” shifted from $\bm{p}_{0}$ or still “resembles” it. While the current population $\bm{p}$ is unknown in practice, $\widehat{\bm{p}}$ serves as an unbiased estimator. We assume that the reference population probabilities satisfy $p_{0j}>0$ for all $j=1,\ldots,B$, ensuring that each category in the reference population has a non-zero probability of occurrence. Additionally, the estimated probabilities $\hat{p}_{j}$, derived from the observed data, are non-negative ($\hat{p}_{j}\geq 0$) for all $j=1,...,B$. Furthermore, the random variable $n\widehat{\bm{p}}$ follows a multinomial distribution with expectation $\mathrm{E}[n\widehat{\bm{p}}]=n\bm{p}$ and variance-covariance matrix $\mathrm{Var}[n\widehat{\bm{p}}]=n\big{[}\mathrm{diag}(\bm{p})-\bm{p}\bm{p}^{\top}\big{]}.$ Here, $\mathrm{diag}(\bm{p})$ is a diagonal matrix with the elements of $\bm{p}$ on the diagonal.
@@ 2.2 Existing measures of population resemblance
Several measures have been proposed for quantifying population shift in credit risk modeling. The most widely used is the Population Stability Index (PSI), introduced by Lewis, (1994). Defined as
@@
This $J$, first introduced by Jeffreys, (1948), is a symmetrized version that addresses the inherent asymmetry of the original “expected per observation information,” here equivalent to $I(\bm{p},\bm{p}_{0})=\sum_{j=1}^{B}p_{j}(\log{{p}_{j}}-\log{{p}_{0j}})$, see Pichler and Schlotter, (2020).
Other measures in risk modeling, inheriting their properties from the Kullback-Leibler diverge, include the Information Value (Siddiqi, , 2017, p. 184) and the Characteristic Stability Index (Siddiqi, , 2017, p. 369). Like the PSI, these measures rely on arbitrary thresholds that disregard sample size and statistical properties.
@@
$\chi^{2}:=\chi^{2}(\bm{p},\bm{p}_{0})=\sum_{j=1}^{B}\frac{(p_{j}-p_{0j})^{2}}{p_{0j}},$
offers a statistically principled alternative for measuring population differences. Notably, both $J$ and $\chi^{2}$ belong to the family of $f$-divergences (Rényi, , 1961; Csiszár, , 1967), with $\chi^{2}$ providing a local approximation to $J$ (and other $f$-divergences) when populations are similar (Csiszár and Shields, , 2004). This property, combined with its well-understood statistical properties, makes $\chi^{2}$ an attractive foundation for population monitoring.
Building on these advantages, we propose the Population Resemblance Statistic (PRS), defined as
@@
offering a measure grounded in well-established statistical theory, thereby also addressing key limitations of the PSI. Specifically, we will develop methods to incorporate sample-size-dependent critical values, using the limiting non-central $\chi^{2}$ properties of the PRS to provide a principled framework for evaluating population shifts.
Other measures include the Kolmogorov-Smirnov (KS) statistic (D’Agostino and Stephens, , 1986), $\mathrm{KS}=\max_{j=1,...,B}\big{|}\hat{F}(j)-F_{0}(j)\big{|}$, where $\hat{F}(j)=\sum_{i=1}^{j}\hat{p}_{i}$ is the empirical cdf and $F_{0}(j)=\sum_{i=1}^{j}p_{0i}$ is the cdf of the model construction population. However, the KS statistic has limited utility for discrete distributions, as its (asymptotic) distribution – unlike in the continuous setting – depends on the underlying $\bm{p}_{0}$ (Conover, , 1972), making critical values less straightforward to determine. The R package dgof offers an implementation of the KS test tailored for the discrete settings (Arnold and Emerson, , 2011). The more recent statistic proposed by Du Pisanie and Visagie, (2020), $\mathrm{DPV}=\max_{j=1,\ldots,B^{\ast}}|\hat{p}_{j}-p_{0j}|/p_{0j}$, relies on an arbitrary selection of $B^{\ast}<B$ and calibration through Monte Carlo methods. Its use has been explored only in settings with $n\geq 10,000$. For further reading on the discrete goodness-of-fit problem in general, see Agresti, (2012).
@@ 2.3 Population resemblance framework
The methodology in this paper relies on the concept of $\delta$-resemblance, as well as additional technical conditions to ensure accurate implementation. Specifically, we assume that the reference probabilities are strictly positive ($p_{0j}>0$ for all $j=1,\ldots,B$) to ensure that every category in the reference distribution is represented. Furthermore, we impose the constraint $0<\delta\leq\min\limits_{j=1,\ldots,B}p_{0j}$, which guarantees that the parameter $\delta$ does not exceed the smallest reference probabilities. This condition effectively limits the maximum allowable shift for any single category to its own probability mass. Finally, $\delta$ is a pre-specified parameter reflecting the organization’s risk tolerance (i.e., the acceptable deviation from the current model), chosen to ensure consistency in decision-making rather than being estimated from the data. However, it is designed to scale with sample size, adapting to variability across different data sets.
The formulation of Definition 1 leverages the Chebyshev distance to quantify the maximum deviation in any single category’s probability. The Chebyshev distance is defined as the largest absolute difference between corresponding elements of two vectors. This distance is particularly advantageous in risk-sensitive applications due to its direct interpretability. By bounding the maximum allowable deviation in any category, the Chebyshev distance provides a precise and actionable metric for monitoring changes in probability distributions.
@@
The Chebyshev distance aligns with the operational need to control critical deviations, as it isolates the largest shift in the data. This characteristic is especially important in credit risk monitoring, where changes in high-risk categories may have outsized implications. Unlike alternatives such as the Euclidean distance, which averages deviations across categories, the Chebyshev distance focuses on the worst-case deviation, offering sharper insights into localized shifts.
@@
In a credit risk management application, consider the portfolio-level probability of default (PD), computed as $PD(\bm{p})=\sum_{i=1}^{B}n_{i}p_{i}$, where $n_{i}$ is the number of borrowers in risk grade $i$, and $p_{i}$ represents the default rate for that grade. When two populations, $\bm{p}$ and $\bm{p}’$ ’, are $\delta$-resemblant, the difference in their portfolio-level PD is bounded by $|PD(\bm{p})-PD(\bm{p}’)|\leq\delta\sum_{i=1}^{B}n_{i}.$ This result demonstrates that the $\delta$-resemblance framework combines a rigorous statistical foundation with practical relevance, enabling risk managers to quantify the population shift’s impact on portfolio performance, assess acceptable levels of deviation aligned with business risk tolerance, and decide when model recalibration is warranted.
@@
We also note that a two-sample formulation of the problem, i.e. treating $\bm{p}_{0}$ as obtained through random sampling from the model construction population, may have some appeal. In practice, however, this approach faces several challenges including dependence between the model construction and current sample, which may stem from at least temporal evolution and overlapping data. To address the issue of non-independence between the samples, we have thus resorted to a one-sample problem formulation, treating the model construction probabilities $\bm{p}_{0}$ as fixed and known. Even in situations as described where the model construction probabilities were established using sampling tools, the one-sample approach can still be employed, providing a conditional inference solution.
The statistical properties of the PSI and the proposed PRS under this framework are explored in Section 3, encompassing both the case of population equality, $\bm{p}=\bm{p}_{0}$, and the more flexible scenario of $\delta$-resemblance, $\bm{p}\in\mathcal{P}(\delta|\bm{p}_{0})$.

### v4

@@ Abstract
keywords:
credit model risk; discrete goodness-of-fit; non-central chi-square; population stability index (PSI); model validation and monitoring; Kullback-Leibler divergence.
@@ 1 Introduction
Testing the stability of a population used for model development is common practice in model risk management. In credit risk modeling, the Population Stability Index (PSI) is the most widely used measure to monitor the evolution of the population underlying a model, through assessing the degree of discrepancy, conversely similarity, between two discrete probability distributions, see [thomas2002credit, pp. 155 ff.] and [siddiqi2017intelligent, pp. 368 ff.]. Small deviations in the population can result in inaccurate or unreliable model predictions. For instance, when modeling the probability of default (or related quantities such as risk score, exposure at default, or loss given default) based on a given population of borrowers, the reliability of the model becomes questionable when the population changes substantively.
Unsurprisingly, many prudential authorities require that the current population resemble (emphasis ours) the one used during regulatory capital model development (see [ecb2024internalmodels, ecb2019validation], [federalreserve2023sr1107a1], [resbank2022credit]; and [pruitt2010applied] for an application of the PSI in SAS). Other areas of application include insurance, healthcare, engineering, and marketing (see [huang], [li], [sahu], [dong2022prediction], [wu2010enterprise], [mcadams2022risk], [chou2022expert], [karakoulas2004empirical] and [brockett1995information]).
@@
Despite its widespread use, the origins and properties of the PSI are not widely understood. The PSI is based on the Kullback-Leibler divergence, measuring difference between two probability distributions [kullback1951information, Eq. 2.6]. The earliest reference to the PSI measure can be found in [lewis1994introduction], who also coined the term “Population Stability Index” and popularized use of the so-called Lewis constants as thresholds (see [thomas2002credit, p. 155 ff.], [siddiqi2017intelligent, p. 368 ff.]). [lewis1994introduction, p. 106] describes the PSI without formulating a hypothesis in the statistical sense but notes, “If a user finds the distribution of scores close together, [they] can be confident that the population has not changed.” In his example, a PSI below 0.10 indicates that the current population is sufficiently similar to the original and no action is required, a value between 0.10 and 0.25 suggests that call for further investigation, and a value above 0.25 signals a material change in the incoming population that may necessitate model reconstruction. The specification of these constants neglects any connection to a defined magnitude of population shift or to the role of sampling variability. By contrast, our framework introduces a precise definition of distributional similarity, which we refer to as resemblance, and explicitly links this to statistical thresholds that specify an associated maximum tolerable population shift.
@@
The arbitrary nature of the Lewis constants has been acknowledged by authors and practitioners alike; see [yurdakul2020statistical], [dupisanievisagie2020], and [moodys2021email]. These thresholds pose limitations in small portfolios, where the shift of a single borrower results in a distortion of the PSI beyond its thresholds and unnecessarily prompts model reconstruction [^1] See [nedbankp3, p. 71], [stbankp3, p. 51] and [frp3, p. 232] for the prevalence of portfolios with less than (e.g.) 100 borrowers.. Conversely, in large portfolios (e.g., retail books with over one million borrowers), substantial shifts may go undetected. Together, these examples highlight the PSI’s limitations as a one-size-fits-all discrepancy measure and highlight the need for more adaptable and interpretable approaches.
1.1 Contributions of this paper
This paper introduces the Population Resemblance Statistic (PRS), a new procedure for monitoring changes in categorical population proportions. Rooted in the Pearson chi-square framework and drawing on properties of the non-central chi-square distribution, the PRS is centered on the concept of $\delta$-resemblance — a formalization of tolerable population shift that allows practitioners to tune the procedure to a specified level of risk tolerance. The resulting critical values for decision-making are directly determined by the sample size and the number of risk categories, making the approach adaptable across a range of settings.
Importantly, the PRS is not merely a marginal refinement of existing tools but represents a conceptual shift in the monitoring paradigm. We move away from the traditional approach, which applies two significance levels (or, in some cases, arbitrary thresholds) to a test statistic under a single point null hypothesis assuming exact population equality. This traditional formulation leads to two well-known problems: in large samples, even negligible shifts are flagged as significant; in small samples, substantial changes may go undetected due to low statistical power – see [schuirmann1987comparison].
By contrast, the PRS approach replaces the point null with a composite null hypothesis. Rather than asking whether any change has occurred, it asks whether the change exceeds a pre-specified, tolerable threshold. This distinction matters. It separates practical and statistical significance, grounding the test in business relevance rather than arbitrary $p$-values. It also enables clearer communication with non-technical stakeholders, as decision thresholds are now expressed in intuitive terms, (e.g., “a shift of more than 5% in any risk category.”)
Finally, the PRS framework supports a more nuanced, three-tiered decision process. By using nested composite hypotheses, it formalizes the familiar “red–amber–green” (RAG) structure in a way that is statistically coherent, flexible, and easily aligned with institutional risk appetite. We also emphasize that the specification of a tolerable shift is not unconstrained: its scale must respect structural limits imposed by the sample size and number of categories. These constraints, and a practical tool for calibration, are presented in Section 3.4.
@@ 1.2 Research aims and objectives
In light of the methodological contributions outlined above, this paper sets out to establish the PRS as a principled and practical alternative to the PSI for monitoring distributional shifts. The PRS framework incorporates explicit risk tolerances, adapts to sample size, and supports decision-making through nested composite hypotheses. The key objectives of this research are as follows:
@@
  • To introduce $\delta$-resemblance as a practical, interpretable way to bound population shift and inform business decisions.
  • To develop the PRS as an alternative monitoring tool with well-calibrated critical values.
  • To formulate the PRS under a composite null hypothesis, relaxing the rigidity of a point null, and justify decision thresholds via the least favorable non-central chi-square distribution.
@@
  • To evaluate PRS performance through simulation studies and real-world applications, highlighting its competency in both small and large sample sizes.
This remainder of the paper is structured as follows: Section 2 formalizes the problem, introducing the PRS as an alternative discrepancy measure with a brief discussion of related methods. Section 3 outlines the statistical properties of the PRS, while Section 3.3 derives the sample-size dependent decision boundaries (critical values). Section 3.4 contains a practical guide to implementing the PRS, including a flowchart for application. Section 4 presents a comprehensive simulation study, and Section 5 applies the PRS to real-world data, comparing it with the widely used PSI and discrete Kolmogorov-Smirnov (KS) statistic. Finally, Section 6 summarizes key contributions and suggests directions for future research.
@@ 2.1 Statistical representation
Consider an independent and identically distributed (i.i.d.) sample of ordinal scores, $X_{1},X_{2},\ldots,X_{n}$, from a population with cumulative distribution function (cdf) $F$ defined on the set of integers $\{1,2,\ldots,B\}$. These scores often arise by discretizing numerical values with each score representing membership in one of $B\geq 2$ disjoint categories (e.g. level of risk). In practice, these categories do not only reflect levels or risk, but are often also determined by operational considerations such as Basel-compliant rating grades, product segmentation or business lines, rather than through formal statistical procedures. As such, the number of categories $B$ are considered fixed a priori.
Let $n_{i}$ denote the count of scores in category $i$, formally expressed as $n_{i}=\sum_{j=1}^{n}\mathbb{I}(X_{j}=i)$, where $\mathbb{I}(A)$ is the indicator function such that $\mathbb{I}(A)=1$ when $A$ is true and $\mathbb{I}(A)=0$ otherwise. The total number of observed scores, $n$, is equal to the sum of all category counts, $n=\sum_{i=1}^{B}n_{i}$.
The true probability of a score falling into category $i$ is $p_{i}=P(X_{j}=i)=F(i)-F(i-1)$ for $i=1,\ldots,B$. Let $\bm{p}=(p_{1},p_{2},\ldots,p_{B})^{\top}$. The observed category proportions, serving as unbiased estimators of the true probabilities, are denoted by $\widehat{\bm{p}}=(\hat{p}_{1},\hat{p}_{2},\ldots,\hat{p}_{B})^{\top}$ with $\hat{p}_{i}=n_{i}/n$ for $i=1,2,\ldots,B$.
The aim is to determine whether the current population $\bm{p}$ is sufficiently close to the reference population $\bm{p}_{0}=(p_{01},p_{02},\ldots,p_{0B})^{\top}$ on which the model was built. The specific nature of this model is not germane. Rather, the question is whether $\bm{p}$ has “substantively” shifted from $\bm{p}_{0}$ or still “resembles” it. While the current population $\bm{p}$ is unknown in practice, $\widehat{\bm{p}}$ serves as an unbiased estimator. We assume that the reference population probabilities satisfy $p_{0i}>0$ for all $i=1,\ldots,B$, ensuring that each category in the reference population has a non-zero probability of occurrence. Additionally, the estimated probabilities, derived from the observed data, are non-negative, i.e. $\hat{p}_{i}\geq 0$ for all i=1,,Bi=1,...,B. Furthermore, the random variable $n\widehat{\bm{p}}$ follows a multinomial distribution with expectation $\mathrm{E}[n\widehat{\bm{p}}\,]=n\bm{p}$ and covariance matrix $\mathrm{Var}[n\widehat{\bm{p}}\,]=n\big[\mathrm{diag}(\bm{p})-\bm{p}\bm{p}^{\top}\big].$ Here, $\mathrm{diag}(\bm{p})$ is a diagonal matrix with the elements of $\bm{p}$ on the diagonal.
@@ 2.2 Existing measures of population resemblance
Several measures have been proposed for quantifying population shift in credit risk modeling. The most widely used is the PSI, introduced by [lewis1994introduction]. Defined as
@@
This $J$, first introduced by [jeffreys1948theory], is a symmetrized version that addresses the inherent asymmetry of the original “expected per observation information,” here equivalent to $I(\bm{p},\bm{p}_{0})=\sum_{j=1}^{B}p_{j}(\log{{p}_{j}}-\log{{p}_{0j}})$, see [pichler2020entropy].
Other measures in risk modeling, inheriting their properties from the Kullback-Leibler diverge, include the Information Value [siddiqi2017intelligent, p. 184] and the Characteristic Stability Index [siddiqi2017intelligent, p. 369]. Like the PSI, these measures rely on arbitrary thresholds that disregard sample size and statistical properties.
@@
$\chi^{2}:=\chi^{2}(\bm{p},\bm{p}_{0})=\sum_{j=1}^{B}\frac{(p_{j}-p_{0j})^{2}}{p_{0j}},$ (3)
offers a statistically principled alternative for measuring population differences. Both $J$ and $\chi^{2}$ are members of the $f$-divergence family [renyi1961measures, csiszar1967information], with $\chi^{2}$ providing a local approximation to $J$ (and other $f$-divergences) when populations are similar [csiszar2004information]. This property, combined with its well-understood statistical behavior, makes $\chi^{2}$ an attractive foundation for population monitoring.
Building on these advantages, we propose the PRS, defined as
@@
which offers a measure grounded in well-established statistical theory and extensively studied properties under a broad range of conditions, thereby also addressing key limitations of the PSI. Specifically, we will develop methods to incorporate sample-size-dependent critical values, using the limiting non-central $\chi^{2}$ properties of the normed PRS to provide a principled framework for evaluating population shifts.
Other measures include the Kolmogorov-Smirnov statistic [dagostino1986goodness], $\mathrm{KS}=\max_{j=1,...,B}\big|\hat{F}(j)-F_{0}(j)\big|$, where $\hat{F}(j)=\sum_{i=1}^{j}\hat{p}_{i}$ is the empirical cdf and $F_{0}(j)=\sum_{i=1}^{j}p_{0i}$ is the cdf of the model construction population. However, the KS statistic has limited utility for discrete distributions, as its (asymptotic) distribution — unlike in the continuous setting — depends on the underlying $\bm{p}_{0}$ [conover1972kolmogorov], making critical values less straightforward to determine. The R package dgof offers an implementation of the KS test tailored for discrete settings [RDGOF]. The more recent statistic of [dupisanievisagie2020], $\mathrm{DPV}=\max_{j=1,\ldots,B^{\ast}}|\hat{p}_{j}-p_{0j}|/p_{0j}$, relies on an arbitrary selection of $B^{\ast}<B$ and Monte Carlo calibration. Its use has been explored only in settings with $n\geq 10{,}000$. For further reading on the discrete goodness-of-fit problem, see [agresti2012categorical].
@@ 2.3 Population resemblance framework
This paper’s methodology builds on $\delta$-resemblance and supporting technical conditions to ensure accurate application. Specifically, we assume that the reference probabilities are strictly positive ($p_{0j}>0$ for all $j=1,\ldots,B$) to ensure that every category in the reference distribution is represented. Furthermore, we impose the constraint $0<\delta\leq\min\limits_{j=1,\ldots,B}p_{0j}$, which guarantees that the parameter $\delta$ does not exceed the smallest reference probabilities. This condition effectively limits the maximum allowable shift for any single category to its own probability mass. Finally, $\delta$ is designed to scale with both the sample size $n$ and the number of categories $B$, allowing it to adapt to variability across different portfolio structures and data conditions. For a given implementation, it is a pre-specified parameter, not estimated from the data, reflecting the organization’s risk tolerance (i.e., the maximum acceptable deviation from the reference population and ensures consistency in decision-making).
@@
The formulation of Definition 1 leverages the Chebyshev distance (also known as the $\ell_{\infty}$ norm) to define a region of tolerable shift, by bounding the largest absolute difference between corresponding category probabilities. This distance is particularly advantageous in risk-sensitive applications due to its direct interpretability and provides a precise, actionable metric for monitoring changes in probability distributions. The Chebyshev distance aligns with the operational need to control critical deviations, as it isolates the largest shift in the data. This characteristic is especially important in credit risk monitoring, where changes in high-risk categories may have outsized implications. Unlike alternatives such as the Euclidean distance, which averages deviations across categories, the Chebyshev distance focuses on the worst-case deviation, offering sharper insights into localized shifts.
@@
In a credit risk management application, consider the portfolio-level probability of default (PD), computed as $\mathrm{PD}(\bm{{p}})=\left(\sum_{i=1}^{B}n_{i}\,\mathrm{pd}_{i}\right)/\sum_{i=1}^{B}n_{i}=\sum_{i=1}^{B}(n_{i}/n)\,\mathrm{pd}_{i}=\sum_{i=1}^{B}p_{i}\,\mathrm{pd}_{i}$ where $n_{i}$ denotes the number of borrowers and $\mathrm{pd}_{i}$ the default probability in each risk category $i$. For two $\delta$-resemblant populations $\bm{p}$ and $\bm{p^{\prime}}$, the difference in their portfolio-level PD is bounded by $|\mathrm{PD}(\bm{p})-\mathrm{PD}(\bm{p^{\prime}})|\leq\delta\sum_{i=1}^{B}\mathrm{pd}_{i}$. This result shows that the $\delta$-resemblance framework balances statistical rigor with practical relevance, enabling risk managers to quantify the population shift’s impact on portfolio performance, assess acceptable levels of deviation aligned with business risk tolerance, and decide when model recalibration is warranted.
@@
We note that a two-sample formulation of the problem, i.e. treating $\bm{p}_{0}$ as arising from a random sample from the model construction population, may have some appeal. In practice, however, this approach is complicated by dependencies between the model construction and current samples, due to temporal evolution or data overlap. To address the issue of non-independence between the populations $\bm{p}_{0}$ and $\bm{p}$, we have thus resorted to a one-sample problem formulation, treating the model construction probabilities $\bm{p}_{0}$ as fixed and known. Even in situations as described where the model construction probabilities were established using sampling tools, the one-sample approach can still be employed, providing a conditional inference solution.